Circular No. 8945

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 J6 (HILL) R. E. Hill reports his discovery of a comet with a diffuse coma but

a clear nuclear condensation and a fan-shaped tail 1′ long in p.a. 210^{o} on four co-added 30-s Catalina Sky Survey exposures obtained in 2′′-4′′ seeing with the 0.68-m Schmidt telescope (discovery observation tabulated below). Following posting on the ‘NEOCP’ webpage, J. C. Pelle writes that images taken by N. Teamo (0.41-m reflector at Punaauia, French Polynesia) on May 14.6 UT show a diffuse coma of diameter 45′′ and a bright central condensation. S. Nakano reports that CCD images taken by Y. Ikari (Moriyama, Shiga-ken, Japan, 0.25-m reflector) on May 14.8 show a coma of diameter 15′′ and a short fan tail in p.a. 240^{o}. E. Guido and G. Sostero write that co-added unfiltered exposures obtained remotely with a 0.25-m reflector near Mayhill, NM, U.S.A., on May 15.4 show a sharp central condensation ∼ 8′′ in diameter, surrounded by a faint, extended coma ∼ 25′′ in diameter (total mag 15.5) and a fan-shaped tail 30′′ long toward the southwest. J. E. McGaha (Tucson, AZ, U.S.A., 0.36-m reflector) reports that his four stacked 60-s exposures with on May 15.4 show a round coma 14′′ in diameter with a tail 48′′ long in p.a. 242^{o}.

2008 UT  α_{2000}  δ_{2000}  Mag.
May 14.46025  22^{h}43^{m}56.^{s}00  +10^{o}47′48′′.4  15.5

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2008-J78.

T = 2008 Apr. 22.838 TT  ω = 16.^{o}385
Ω = 299.274 }2000.0
q = 2.05912 AU  i = 45.904

COMET C/2007 W1 (BOATTINI) Additional visual total-magnitude and coma-diameter estimates (cf.

IAUC 8899, 8931): Apr. 13.82 UT, 8.6, 4′.2 (W. Hasubick, Buchloe, Germany, 10-cm refractor); 23.88, 7.6, 15′ (J. J. Gonzalez, Leon, Spain, 10×50 binoculars); 30.93, 7.3, 12′ (A. Amorim, Florianopolis, Brazil, 10×50 binoculars); May 2.44, 6.9, 18′ (S. Rae, Tauranga, New Zealand, 10×50 binoculars); 6.00, 7.0, 25′ (M. Goiato, Araçatuba, Brazil, 7×50 binoculars); 8.45, 6.4, – (D. A. J. Seargent, Cowra, NSW, Australia, naked eye); 12.02, 6.5, 15′ (Amorim; moonlight); 14.96, 6.2, 23′ (Goiato; moonlight).

2008 May 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green